OpenStudy (jessiegonzales):

MEDA!!! Use an algebraic equation to solve the problem. A rectangle is 3 times as long as it is wide. The perimeter is 60cm. Find the dimension of the rectangle. Round to the nearest tenth if necessary. A. 7.5cm by 22.5cm B. 7.5cm by 52.5cm C. 20cm by 60cm D. 15cm by 22.5cm

OpenStudy (jacob902):

We know the perimeter of a rectangle is 2 * width + 2 * length, and we know that the length is 3* width, so: 2W + 2L = 60 L = 3W Now we can substitute 3W for L: 2W + 2L = 60 2W + 2(3W) = 60 2W + 6W = 60 8W = 60 W = 7.5 L = 3W = 3(7.5) = 22.5 So the rectangle is 7.5 x 22.5